OpenStudy (anonymous):

If a/b = c/d = k and a,b,c,d are positive, then which of the following must be true? A. (a+c)/(b+d)=k B. ab=cd=k C. ac=bd=k D. a=c=k E. ac/bd = k

hartnn (hartnn):

if a/b = c/d can i tell, a/c = b/d ?

hartnn (hartnn):

what i did is multiply b on both sides and divide both sides by c

OpenStudy (anonymous):

times b/c right?

hartnn (hartnn):

yup

hartnn (hartnn):

a/c = b/d now add 1 on both sides, what u get ?

OpenStudy (anonymous):

add 1??

hartnn (hartnn):

yup, add 1 to both sides and see what happens ? a/c +1 = ... ? b/d+1 = ...?

OpenStudy (anonymous):

a+c / c b+d / d

hartnn (hartnn):

correct so, (a+c)/c = (b+d)/d can you rearrange this ?

OpenStudy (anonymous):

a+c / b+d = c/d = k so A is correct

hartnn (hartnn):

from that what does c/d equal ??

OpenStudy (anonymous):

c/d=k

hartnn (hartnn):

yup, you got it :)

OpenStudy (anonymous):

so A must be true?

hartnn (hartnn):

correct! :)

OpenStudy (anonymous):

why B,C,D,E are not true then?

hartnn (hartnn):

because ab, won't equal 'k' ac = bd, but this won't equal k again, ac = bd = kbd a and c would not = k, unless b,d =1 lastly, a/c =k, b/d = k, ac/bd = k*k = k^2 , not k

OpenStudy (anonymous):

ok, i got it thanks:)

hartnn (hartnn):

welcome ^_^